# SKADA-Bench: Benchmarking Unsupervised Domain Adaptation Methods with Realistic Validation

### v1

@@ D.4 Comparions between supervised and unsupervised scorers
We observe in Figure D.1 the cross-validation score as a function of the final accuracy for various DA methods type and for both supervised and unsupervised scorers. As expected, we observe a good correlation between accuracy and cross-validation score with the supervised scorer. An important remark is that the Circular Validation (CircV) [4] shows some correlation between accuracy and cross-validation score. It indicates that this unsupervised scorer might be the most suitable choice for hyperparameter selection. This is supported by our extended experimental results in Table 2 for which the CircV is selected as the best scorer the most often. A similar trend can be observed for the Importance Weighted (IW) [49] which is also confirmed in Table 2.
Figure D.1: Cross-val score as a function of the accuracy for various DA methods and different supervised and unsupervised scorers. Each point represents an inner split with a DA method (color of the points) and a dataset. A good scorer should have a score that correlates with the target accuracy.
Figure D.2: Change of accuracy of the DA methods with the best realistic unsupervised scorer (Table 2) w.r.t. the supervised scorer.
@@
We plot the loss in performance of the DA methods with the best realistic unsupervised scorer compared to using the supervised scorers in Figure D.2.
@@
We present a scatter plot in Figure D.3 and Figure D.4 , the accuracy of different DA methods using both supervised scorer and unsupervised scorer. In this figure, points below the diagonal indicate a decrease in performance when using the unsupervised scorer compared to the supervised one. The colors represent different types of DA methods. We can see that the SND, DEV and PE scorers all lead to a large performance loss compared to the supervised scorer. While IW and CircV results are much more concentrated near the diagonal, indicating a small loss in performance. This concentration explains why these two scorers have been selected as the best scorers for most of the methods in Table 2.
Figure D.3: Accuracy of the DA methods using unsupervised scorers as a function of the accuracy with the supervised scorer. Colors represent the type of DA methods.
Figure D.4: Accuracy of the DA methods using unsupervised scorers as a function of the accuracy with the supervised scorer for the different types of DA methods. Points below the diagonal represent a decrease in performance when using the unsupervised scorer compared to the supervised one. Colors represent the dataset on which the DA method is applied.
D.5 Computational efficiency of the DA methods
Figure D.5 shows the average computation time for training and testing each method. These results are based on one outer split, while we ran the benchmarks for five outer splits. Each method has a different time complexity. Interestingly, more time-consuming methods are not necessarily more performant than others. For instance, the highest-ranked methods—LinOT, CORAL, and SA—also have some of the lowest training and testing times. It’s also worth noting that during the experiments, we enforced a 4-hour timeout. Thus, the more time-intensive methods might have been even slower without this timeout.
Figure D.5: Mean computing time to train and test each method for every experiment outer split.
D.6 Hyperparameters grid search for the DA methods
In this section, we report the grids of hyperparameters used in our grid search for each DA method. These values are also available in our code that is provided in the supplementary material.
Table D.13: Hyperparameter grids used in the grid search for each DA method. The hyperparameter grids were designed to be minimal yet expressive, allowing each method to perform optimally. We selected parameters based on what seemed most reasonable, according to our best knowledge.

### v2

@@ E.6 Comparisons between supervised and unsupervised scorers
We observe in Figure 6 the cross-validation score as a function of the final accuracy for various DA methods type and for both supervised and unsupervised scorers. As expected, we observe a good correlation between accuracy and cross-validation score with the supervised scorer. An important remark is that the Circular Validation (CircV) (Bruzzone & Marconcini, 2010b) shows some correlation between accuracy and cross-validation score. It indicates that this unsupervised scorer might be the most suitable choice for hyperparameter selection. This is supported by our extended experimental results in Table 2 for which the CircV is selected as the best scorer the most often. A similar trend can be observed for the Importance Weighted (IW) (Sugiyama et al., 2007a) which is also confirmed in Table 2.
Figure 6: Cross-val score as a function of the accuracy for various DA methods and different supervised and unsupervised scorers. Each point represents an inner split with a DA method (color of the points) and a dataset. A good scorer should have a score that correlates with the target accuracy.
Figure 7: Change of accuracy of the DA methods with the best realistic unsupervised scorer (Table 2) w.r.t. the supervised scorer.
@@
We plot the loss in performance of the DA methods with the best realistic unsupervised scorer compared to using the supervised scorers in Figure 7.
@@
We present a scatter plot in Figure 8 and Figure 9 , the accuracy of different DA methods using both supervised scorer and unsupervised scorer. In this figure, points below the diagonal indicate a decrease in performance when using the unsupervised scorer compared to the supervised one. The colors represent different types of DA methods. We can see that the SND, DEV and PE scorers all lead to a large performance loss compared to the supervised scorer. While IW and CircV results are much more concentrated near the diagonal, indicating a small loss in performance. This concentration explains why these two scorers have been selected as the best scorers for most of the methods in Table 2.
Figure 8: Accuracy of the DA methods using unsupervised scorers as a function of the accuracy with the supervised scorer. Colors represent the type of DA methods.
Figure 9: Accuracy of the DA methods using unsupervised scorers as a function of the accuracy with the supervised scorer for the different types of DA methods. Points below the diagonal represent a decrease in performance when using the unsupervised scorer compared to the supervised one. Colors represent the dataset on which the DA method is applied.
E.7 Computational efficiency of the DA methods
Figure 10 shows the average computation time for training and testing each method. These results are based on one outer split, while we ran the benchmarks for five outer splits. Each method has a different time complexity. Interestingly, more time-consuming methods are not necessarily more performant than others. For instance, the highest-ranked methods—LinOT, CORAL, and SA—also have some of the lowest training and testing times. It’s also worth noting that during the experiments, we enforced a 4-hour timeout. Thus, the more time-intensive methods might have been even slower without this timeout.
Figure 10: Mean computing time to train and test each method for every experiment outer split.
E.8 Hyperparameters grid search for the DA methods and neural networks training
In this section, we first report the grids of hyperparameters used in our grid search for each DA method.
We also detail the configuration and hyperparameter grids for training neural networks in our Deep DA benchmark. We provide an overview of the key settings for each dataset, including batch sizes, optimizer parameters, learning rates, and the number of training epochs. Additionally, we outline the hyperparameter grids used for grid search across the Deep DA methods.
Table 23: Hyperparameter grids used in the grid search for each DA method. The hyperparameter grids were designed to be minimal yet expressive, allowing each method to perform optimally. We selected parameters based on what seemed most reasonable, according to our best knowledge.
@@ E.8 Hyperparameters grid search for the DA methods and neural networks training
Table 24: Configuration of Deep learning models for each dataset. This includes recommended Batch sizes, optimizer settings, learning rates, and maximum epochs.
Dataset Configuration
mnist_usps Neural net: 2-layer CNN Batch size: 256 Optimizer: SGD, momentum=0.6, weight_decay=1e-5 Learning rate: 0.1 Epochs: 20 Learning rate scheduler: LRScheduler(StepLR, step_size=10, gamma=0.2)
office31 Neural net: ResNet50 Batch size: 128 Optimizer: SGD, momentum=0.2, weight_decay=1e-5 Learning rate: 0.5 Epochs: 30 Learning rate scheduler: StepLR, step_size=10, gamma=0.2
officehome Neural net: ResNet50 Batch size: 128 Optimizer: SGD, momentum=0.6, weight_decay=1e-5 Learning rate: 0.05 Epochs: 20 Learning rate scheduler: StepLR, step_size=10, gamma=0.2
bci Neural net: FBCSPNet Batch size: 64 Optimizer: AdamW Learning rate: 0.000625 Epochs: 200 Learning rate scheduler: CosineAnnealingLR
Table 25: Hyperparameter grids used in the grid search for each Deep DA method. The hyperparameter grids were designed to be minimal yet expressive, allowing each method to perform optimally. We selected parameters based on what seemed most reasonable, according to our best knowledge.
Method Hyperparameter Grid
DANN ’reg’: [0.001, 0.01, 0.1, 1.0]
DeepCORAL ’reg’: [0.00001, 0.0001, 0.001, 0.01, 0.1, 1.0, 10.0, 100.0, 1000.0]
DeepJDOT ’reg_cl’: [0.0001, 0.001, 0.01]
’reg_dist’: [0.0001, 0.001, 0.01]
DAN ’reg’: [0.00001, 0.0001, 0.001, 0.01, 0.1, 1.0, 10.0, 100.0, 1000.0]
MCC ’reg’: [0.01, 0.1, 1]
’temperature’: [1, 2, 3]
MDD ’reg’: [0.001, 0.01, 0.1]
’gamma’: [1, 3]
SPA ’reg’: [0.001, 0.01, 0.1, 1]
’reg_nap’: [0, 1]